# EnKF-C user guide version 2.32.0

### v11

@@ Introduction
Following is the list of main pre-requisites and limitations resulted from the design and algorithmic solutions adopted in EnKF-C:
@@ 1.1 Kalman filter
The Kalman filter (KF) is the underlying concept behind the EnKF. It is rather simple if formulated as the recursive least squares.
@@
This solution is known as the Kalman filter (KF, Kalman, 1960). Equations (1.7) describe advancing the system in time and represent the stage commonly called “forecast”, while equations (1.6) describe assimilation of observations and represent the stage called “analysis”. The superscripts $f$ and $a$ are used hereafter to refer to the forecast and analysis variables, correspondingly. The forecast and analysis model state estimates $\mathbf{x}^{f}$ and $\mathbf{x}^{a}$ are commonly called (simply) forecast and analysis. Matrix $\mathbf{K}$ is called Kalman gain.
@@ 1.2 EnKF
The standard form of the KF (1.7, 1.6) is not necessarily the most convenient or suitable one in practice. The corresponding algorithms can be prone to loosing the positive definiteness of the state error covariance $\mathbf{P}$ due to round-up errors; and more importantly, explicit use of $\mathbf{P}$ makes these algorithms non-scalable in regard to the model state dimension.
Both these immediate problems can be addressed with the ensemble Kalman filter, or the EnKF. In the EnKF the SDAS is carried by an ensemble of $m$ model states $\mathbf{E}$, which can be split into ensemble mean and ensemble anomalies:
@@
[…]
@@
[…]
@@ 1.3 EnKF analysis
In this section we will give a brief overview of solutions for the EnKF analysis, and then describe the particular schemes used in EnKF-C.
@@ 1.3.2 Some schemes
Correction. The Andrews’ solution turns out to be right only for diagonal $\mathbf{R}$. The correct solution of form (1.28) can be obtained from (1.27) by using identity $\mathbf{M}^{-1/2}=\mathbf{I}-(\mathbf{M}+\mathbf{M}^{1/2})^{-1}(\mathbf{M}-\mathbf{I})$:
@@
where $\mathbf{M}_{m}\equiv\mathbf{I}+\mathbf{S}^{\tiny\mathrm{T}}\mathbf{S}$, $\mathbf{M}_{p}\equiv\mathbf{I}+\mathbf{S}\mathbf{S}^{\tiny\mathrm{T}}$, $\mathbf{S}\equiv\mathbf{R}^{-1/2}\mathbf{H}\mathbf{A}^{f}/(m-1)^{1/2}$ (boc16a).
@@
where $\hat{\mathbf{M}}^{1/2}\equiv\mathbf{R}^{1/2}(\mathbf{R}^{-1/2}\mathbf{M}\mathbf{R}^{-1/2})^{1/2}$. (1.30) concides with the Andrews’ solution (1.28) if $\hat{\mathbf{M}}^{1/2}=\mathbf{M}^{1/2}$, which is the case for diagonal $\mathbf{R}$.
@@ 1.3.4 Ensemble reduction
where $\mathcal{F}(\cdot)$ is some function and $\mathbf{S}$ are the standardised ensemble observations (1.34), then the analysed ensemble can be downsized as follows:
@@ 1.4 Localisation
There are two common localisation methods for the EnKF – covariance localisation (CL, Hamill and Whitaker, 2001; Houtekamer and Mitchell, 2001), also known as covariance filtering, and local analysis (LA, Evensen, 2003; Ott et al., 2003, rev. 2005). Although CL may have advantages in certain situations (non-local observations, “strong” assimilation), in practice the two methods produce similar results (Sakov and Bertino, 2011). For algorithmic reasons EnKF-C uses LA.
[…]
@@ 1.5 Asynchronous DA
Observations assimilated at each cycle in the KF are assumed to be made simultaneously at the time of assimilation. In such cases observations and DA method are referred to as synchronous. In reality, observations assimilated at a given cycle are made over some period of time called “data assimilation window” (DAW). If the DA method accounts for the time of observations, observations and DA method are referred to as asynchronous.
@@
[…]
@@ 2.2 The workflow
EnKF-C conducts data assimilation in three stages: PREP, CALC and UPDATE.
PREP preprocesses observations so that they are ready for DA. It has the following stages:
@@
PREP does not need to access the model state; it only needs to access the model grid. (Note that for some types of vertical coordinates the vertical model grid depends on the state.)
@@
Apart from this main mode of operation, CALC can also be used for calculating the forecast innovations, or operate in the single observation experiment mode.
UPDATE updates the ensemble (EnKF) or the background (EnOI) using the transforms calculated by CALC, along with a number of specified diagnostics, such as the ensemble spread, inflation, or vertical correlations of subsurface fields with the surface field.
The principle diagram of EnKF-C workflow is shown in Fig. 2.1.
Figure 2.1: The principle diagram of EnKF-C workflow.
@@ 2.3 Starting up: example 1
It may be a good idea to start getting familiar with the system by running the example in examples/1. The example has been put up based on runs of the regional EnKF and EnOI reanalysis systems for Tasman Sea developed by Bureau of Meteorology. It allows one to conduct a single assimilation for 23 December 2007 (day 6565 since 1 January 1990) with either EnKF or EnOI. To reduce the size of the system, the model state has been stripped down to two vertical levels and $100\times 100$ horizontal grid. Due to its size (almost 80 MB) the data for this example is available for download separately from the EnKF-C code – see examples/1/README for details.
@@ 2.4.1 Main parameter file
The main parameter file specifies the main parameters of DA and 4 other parameter files. Its format is described by running enkf_prep, enkf_calc or enkf_update with option --describe-prm-format:
@@
     >./bin/enkf_prep --describe-prm-format      Main parameter file format:        TIME            = {<# days since YYYY-MM-DD> | <value>}    [ WINDOWMIN       = <start of obs window in days from analysis> ] (-inf*)    [ WINDOWMAX       = <end of obs window in days from analysis> ]   (+inf*)      MODE            = { ENKF | ENOI | HYBRID }    [ SCHEME          = { DENKF* | ETKF } ]                    (MODE = ENKF or HYBRID)    [ ALPHA           = <alpha> ]                              (1*) (MODE = ENKF or HYBRID)      GAMMA           = <gamma>                                (MODE = HYBRID)      MODEL           = <model prm file>      GRID            = <grid prm file>      OBSTYPES        = <obs. types prm file>      OBS             = <obs. data prm file>      ENSDIR          = <ensemble directory>                   (except MODE = ENOI and                                                               --forecast-stats-only)    [ ENSDIR_STATIC   = <static ensemble directory> ]          (MODE = HYBRID)    [ ENSSIZE         = <total ensemble size> ]                (<full>*)    [ ENSSIZE_DYNAMIC = <size of dynamic ensemble> ]           (<full>*) (MODE = HYBRID)    [ ENSSIZE_STATIC  = <size of static ensemble> ]            (<full>*) (MODE = HYBRID)      BGDIR           = <background directory>                 (MODE = ENOI)    [ KFACTOR         = <kfactor> ]                            (NaN*)    [ RFACTOR         = <rfactor> ]                            (1*)      LOCRAD          = <loc. radius in km> ...      LOCWEIGHT       = <loc. weight> ...                      (# LOCRAD > 1)    [ NLOBSMAX        = <max. number of local obs. of each type> ]    [ STRIDE          = <stride for ensemble transforms> ]     (1*)    [ SOBSTRIDE       = <stride for superobing> ]              (1*)    [ FIELDBUFFERSIZE = <fieldbuffersize> ]                    (1*)    [ INFLATION       = <inflation> [ <VALUE>* | PLAIN ]       (1*)    [ REGION          = <name> <lon1> <lon2> <lat1> <lat2>      ...    [ POINTLOG        <lon> <lat> [grid name]]      ...    [ EXITACTION      = { BACKTRACE* | SEGFAULT } ]    [ BADBATCHES      = <obstype> <max. bias> <max. mad> <min # obs.> ]      ...    [ ZSTATINTS       = [<z1> <z2>] ... ]    [ NCFORMAT        = { CLASSIC | 64BIT | NETCDF4 } ]        (NETCDF4*)    [ NCCOMPRESSION   = <compression level> ]                  (0*)      Notes:      1. { ... | ... | ... } denotes the list of possible choices      2. [ ... ] denotes an optional input      3. ( ... ) is a note      4. * denotes the default value      5. < ... > denotes a description of an entry      6. ... denotes repeating the previous item an arbitrary number of times      7. Depending on the context, some of the entries may be redundant      8. TIME entry is also a flag for geophysical/non-geophysical system,         depending on the presence of "days since"     The TIME entry also serves as a switch between geophysical/non-geophysical system. For geophysical systems the coordinates are assumed to be geographic longitude and latitude (unless GEOGRAPHIC = 0 is specified for a grid); the time is assumed to be dimensional, in some physical units; and the localisation radii is assumed to be in kilometers. For non-geophysical systems all the above are assumed to be non-dimensional, and the grid is assumed to be on a plane.
@@ 2.4.2 Model parameter file
The model parameter file mainly describes the composition of the state vector by listing the model variables and specifying the associated grids.
@@
Each model variable is described in a block started by the entry for the variable name. The inflation parameters for a variable, if specified, override the common values set in the main parameter file (sec. 2.8.1). The option APPLYLOG makes it possible to conduct assimilation in log space (sec. 2.15), and RANDOMISE – to specify a “forgetting” model for the variable (sec. 2.14).
EnKF-C permits using multiple model grids. In this case each model variable must be associated with one of the grids defined in the grid parameter file. See examples/4 for an example.
@@ 2.4.3 Grid parameter file
     >./bin/enkf_prep --describe-prm-format grid      Grid parameter file format:        NAME             = <name> [ PREP | CALC ]    [ DOMAIN           = <domain name> ]      DATA             = <data file name>      (either)        HTYPE          = { rect | curv | unstr | none }        XVARNAME       = <X variable name>        YVARNAME       = <Y variable name>        (if htype = unstr)        [ TRIVARNAME    = <triangle vertice IDS variable name> ]        [ TRINEIVARNAME = <triangle neighbour IDS variable name> ]        (end if)      (or)        HGRIDFROM      = <grid name>      (end either)      VTYPE            = { z | sigma | hybrid | numeric | none }    [ VDIR             = { fromsurf* | tosurf } ]    [ GEOGRAPHIC       = { 0 | 1* | 2 } ]      (if vtype = z)        ZVARNAME       = <Z variable name>      [ ZCVARNAME      = <ZC variable name> ]      [ NUMLEVELSVARNAME = <# of levels variable name> ]      [ DEPTHVARNAME   = <depth variable name> ]      (else if vtype = sigma)        CVARNAME       = <Cs_rho variable name>      [ CCVARNAME      = <Cs_w variable name> ]      [ SVARNAME       = <s_rho variable name> ]            (uniform*)      [ SCVARNAME      = <s_w variable name> ]              (uniform*)      [ HCVARNAME      = <hc variable name> ]               (0.0*)      [ DEPTHVARNAME   = <depth variable name> ]      [ MASKVARNAME    = <land mask variable name> ]      (else if vtype = hybrid)        AVARNAME       = <A variable name>        BVARNAME       = <B variable name>      [ ACVARNAME      = <AC variable name> ]      [ BCVARNAME      = <BC variable name> ]        P1VARNAME      = <P1 variable name>        P2VARNAME      = <P2 variable name>      [ MASKVARNAME    = <land mask variable name> ]      (else if vtype = numeric)        ZVARNAME       = <Z variable name>      [ ZCVARNAME      = <ZC variable name> ]      [ NUMLEVELSVARNAME = <land mask variable name> ]      [ DEPTHVARNAME   = <depth variable name> ]       (end if)    [ STRIDE           = <stride for ensemble transforms> ] (1*)    [ SOBSTRIDE        = <stride for superobing> ]          (1*)    [ ZSTATINTS        = [<z1> <z2>] ... ]      [ <more of the above blocks> ]
While the code is supposed to automatically identify the type of horizontal grid used, it can also be specified explicitly. The type of the vertical grid has to be specified. If some grid (say, grid A) has the same horizontal grid as other grid (grid B), the code can be notified of this by entering name of grid A in the field HGRIDFROM for grid B (or vice versa). This saves memory, grid initialisation time, and uses transforms calculated for grid A for all variables defined on grid B.
@@
    rectangular quadrilateral grids aligned with physical coordinates;
@@
The flag GEOGRAPHIC introduced in v2.8.0 if set to “1” internally communicates to the code that the grid nodes are defined in geographic coordinates and sets grid rendering to be performed in 3D cartesian coordinates. This can be essential in polar regions because of the skewness of grid cells in geographic coordinates there (it can cause failures of the mapping algorithm). The user can investigate the necessity of setting this flag to “yes” by comparing the numbers of accepted/rejected observations with and without it.
@@
The flag GEOGRAPHIC set to “2” treats grid coordinates similarly to that for the flag set to “1”, but different rendering procedures are used, which may turn out more robust for e.g. grids with artificial discontinuities in longitude.
For unstructured grids associated with observed variables it is necessary to specify triangulation details: vertice IDs and neighbour triangle IDs. For unstructured grids stride can be equal to 1 only (i.e. ensemble transforms must be calculated in each grid node).
@@
The type of the vertical grid is defined by entry VTYPE in the grid parameter file. EnKF-C supports 3 types of vertical grids: $Z$ (“z”); $\sigma$ (“sigma”); and hybrid $\sigma$-$p$ ($\sigma$ pressure, “hybrid”). It is also possible to define a purely horizontal (two-dimensional) grid by defining its vertical type as “none”.
@@
Vertical grid type numeric is supposed to handle general situations described by 3D arrays of layer depths. Grids of this type are defined by specifying variable with cell centres depth coordinates (entry ZVARNAME) and (optionally) variable with cell boundaries depth coordinates (entry ZCVARNAME).
@@
“Empty” grids have been introduced in version 2.28.0 to simplify estimation of global scalar parameters. In the KF context a model paramater is simply an unobserved model variable. In EnKF-C model variables are defined in the model parameter file; each model variable must be associated with a grid defined in the grid parameter file. “Empty” grids are used to be associated with global model paramaters and are defined by specifying their vertical and horizontal types as “none”, e.g.

### v12

# EnKF-C user guide version 2.43.4

@@ Introduction
Following is the list of main prerequisites and limitations resulting from the design and algorithmic solutions adopted in EnKF-C:
@@ 1.1 Kalman filter
The Kalman filter (KF) is the underlying concept behind the EnKF. It is rather simple if formulated as recursive least squares.
@@
This solution is known as the Kalman filter (KF, Kalman, 1960). Equations (1.7) describe advancing the system in time and represent the stage commonly called “forecast”, while equations (1.6) describe assimilation of observations and represent the stage called “analysis”. The superscripts $f$ and $a$ are used hereafter to refer to the forecast and analysis variables, respectively. The forecast and analysis model state estimates $\mathbf{x}^{f}$ and $\mathbf{x}^{a}$ are commonly called (simply) forecast and analysis. Matrix $\mathbf{K}$ is called Kalman gain.
@@ 1.2 EnKF
The standard form of the KF (1.7, 1.6) is not necessarily the most convenient or suitable one in practice. The corresponding algorithms can be prone to losing the positive definiteness of the state error covariance $\mathbf{P}$ due to rounding errors; and more importantly, explicit use of $\mathbf{P}$ makes these algorithms non-scalable in regard to the model state dimension.
Both these immediate problems can be addressed with the ensemble Kalman filter (EnKF). In the EnKF the SDAS is carried by an ensemble of $m$ model states $\mathbf{E}$, which can be split into ensemble mean and ensemble anomalies:
@@
$\mathcal{H}(\mathbf{x})\quad$ $\to$ $\quad\mathcal{H}(\mathbf{E})\,\mathbf{1}/m$ (1.11b)
@@
$\mathcal{M}(\mathbf{x})\quad$ $\to$ $\quad\mathcal{M}(\mathbf{E})\,\mathbf{1}/m$ (1.11d)
@@ 1.3 EnKF analysis
In this section we will give a brief overview of solutions for EnKF analysis and describe the particular schemes used in EnKF-C.
@@ 1.3.2 Some schemes
Correction. The Andrews’ solution turns out to be valid only for diagonal $\mathbf{R}$. The correct solution of form (1.28) can be obtained from (1.27) by using identity $\mathbf{M}^{-1/2}=\mathbf{I}-(\mathbf{M}+\mathbf{M}^{1/2})^{-1}(\mathbf{M}-\mathbf{I})$:
@@
where $\mathbf{M}_{m}\equiv\mathbf{I}+\mathbf{S}^{\tiny\mathrm{T}}\mathbf{S}$, $\mathbf{M}_{p}\equiv\mathbf{I}+\mathbf{S}\mathbf{S}^{\tiny\mathrm{T}}$, $\mathbf{S}\equiv\mathbf{R}^{-1/2}\mathbf{H}\mathbf{A}^{f}/(m-1)^{1/2}$ (Bocquet, 2016).
@@
where $\hat{\mathbf{M}}^{1/2}\equiv\mathbf{R}^{1/2}(\mathbf{R}^{-1/2}\mathbf{M}\mathbf{R}^{-1/2})^{1/2}$. (1.30) coincides with the Andrews’ solution (1.28) if $\hat{\mathbf{M}}^{1/2}=\mathbf{M}^{1/2}$, which is the case for diagonal $\mathbf{R}$.
@@ 1.3.4 Ensemble reduction
where $\mathcal{F}(\cdot)$ is some function and $\mathbf{S}$ are the standardised ensemble observation anomalies (1.34), then the analysed ensemble can be downsized as follows:
@@ 1.4 Localisation
There are two common localisation methods for the EnKF – covariance localisation (CL, Hamill and Whitaker, 2001; Houtekamer and Mitchell, 2001), also known as covariance filtering, and local analysis (LA, Houtekamer and Mitchell, 1998; Evensen, 2003; Ott et al., 2003, rev. 2005). Although CL may have advantages in certain situations (non-local observations, “strong” assimilation), in practice the two methods produce similar results (Sakov and Bertino, 2011). For algorithmic reasons EnKF-C uses LA.
Instead of calculating the global ensemble transform $\mathbf{X}_{5}$, LA involves calculating local ensemble transforms $\accentset{i}{\mathbf{X}}_{5}$ for each element $i$ of the state vector. This is done using local normalised ensemble observation anomalies $\accentset{i}{\mathbf{S}}$ and local normalised innovation $\accentset{i}{\mathbf{s}}$, obtained by tapering global $\mathbf{S}$ and $\mathbf{s}$:
$\accentset{i}{\mathbf{s}}\equiv\mathbf{s}\circ\accentset{i}{\mathbf{f}},$ (1.42a)
$\accentset{i}{\mathbf{S}}\equiv\mathbf{S}\circ(\accentset{i}{\mathbf{f}}\,\mathbf{1}^{\tiny\mathrm{T}}),$ (1.42b)
where $\accentset{i}{\mathbf{f}}$ is the vector of taper coefficients for element $i$, and $\mathbf{A}\circ\mathbf{B}$ denotes by-element, or Hadamard, or Schur product of matrices $\mathbf{A}$ and $\mathbf{B}$. We consider non-adaptive localisation only, when the taper coefficient is a function of locations of the state element $i$ (denoted as $\accentset{i}{\mathbf{r}}$) and observation $o$ (denoted as $\accentset{\{o\}}{\mathbf{r}}$$)\accentset{i}{\mathbf{f}}_{o}=g(\accentset{i}{\mathbf{r}},\accentset{\{o\}}{\mathbf{r}}$), where $g$ is the taper function. In layered geophysical models $g$ is often assumed to depend only on horizontal distance between these locations:
@@ 1.5 Asynchronous DA
Observations assimilated at each cycle in the KF are assumed to be made simultaneously at the time of assimilation. In such cases, observations and DA method are referred to as synchronous. In reality, observations assimilated at a given cycle are made over some period of time called “data assimilation window” (DAW). If the DA method accounts for the time of observations, observations and DA method are referred to as asynchronous.
@@
Note that usually only $\mathcal{H}$ is a function that depends on assimilated observations; now by introducing $\mathcal{H}\circ\mathcal{M}(\mathbf{x})=\mathcal{H}[\mathcal{M}(\mathbf{x})]$ we have to assume that $\mathcal{M}$ also depends on observations, propagating the initial state to the time of each observation. It is also possible to interpret $\mathcal{M}(\mathbf{x}$) as the trajectory starting from $\mathbf{x}$, while $\mathcal{H}$ maps it to observations.
@@ 2.2 The workflow
EnKF-C conducts data assimilation in three stages: PREP, CALC, and UPDATE, by running executables enkf_prep, enkf_calc, and enkf_update, correspondingly. EnKF-C also provides ens_diag for calculating a number of ensemble diagnostic variables.
PREP preprocesses observations so that they are ready for DA. It has the following steps:
@@
PREP does not need to access the model state; it only accesses the model grid. (Note that for some types of vertical coordinates the vertical model grid may depend on the state.)
@@
Apart from this main mode of operation, CALC can also be used for calculating forecast innovations, or operate in the single observation experiment mode.
UPDATE updates the ensemble (EnKF) or the background (EnOI) using the transforms calculated by CALC, along with a number of specified diagnostics, such as the ensemble spread and inflation.
The principle diagram of EnKF-C workflow in EnKF mode is shown in Fig. 2.1.
Figure 2.1: The principle diagram of EnKF-C workflow in EnKF mode.
@@ 2.3 Starting up: example 1
You may find it helpful to start getting familiar with the system by running the example in examples/1. The example has been put up based on runs of the regional EnKF and EnOI reanalysis systems for Tasman Sea developed by Bureau of Meteorology. It allows one to conduct a single assimilation for 23 December 2007 (day 6565 since 1 January 1990) with either EnKF or EnOI. To reduce the size of the system, the model state has been stripped down to two vertical levels and $100\times 100$ horizontal grid. Due to its size (almost 80 MB) the data for this example is available for download separately from the EnKF-C code – see examples/1/README for details.
@@ 2.4.1 Main parameter file
The main parameter file specifies the main parameters of DA and four other parameter files. Its format is described by running enkf_prep, enkf_calc or enkf_update with option --describe-prm-format:
@@
      >./bin/enkf_prep --describe-prm-format         Main parameter file format:         TIME = {<# days since YYYY-MM-DD> | <value>}     [ WINDOWMIN = <start of obs window in days from analysis> ] (-inf*)     [ WINDOWMAX = <end of obs window in days from analysis> ] (+inf*)     MODE = { ENKF | ENOI | HYBRID }     [ SCHEME = { DENKF* | ETKF } ] (MODE = ENKF or HYBRID)     [ ALPHA = <alpha> ] (1*) (MODE = ENKF or HYBRID)     GAMMA = <gamma> (MODE = HYBRID)     MODEL = <model prm file>     GRID = <grid prm file>     OBSTYPES = <obs. types prm file>     OBS = <obs. data prm file>     ENSDIR = <ensemble directory> (except MODE = ENOI and     --forecast-stats-only)     [ ENSDIR_STATIC = <static ensemble directory> ] (MODE = HYBRID)     [ ENSSIZE = <total ensemble size> ] (<full>*)     [ ENSSIZE_DYNAMIC = <size of dynamic ensemble> ] (<full>*) (MODE = HYBRID)     [ ENSSIZE_STATIC = <size of static ensemble> ] (<full>*) (MODE = HYBRID)     BGDIR = <background directory> (MODE = ENOI)     [ KFACTOR = <kfactor> ] (NaN*)     [ RFACTOR = <rfactor> ] (1*)     LOCRAD = <loc. radius in km> ...     LOCWEIGHT = <loc. weight> ... (# LOCRAD > 1)     [ NLOBSMAX = <max. number of local obs. of each type> ]     [ STRIDE = <stride for ensemble transforms> ] (1*)     [ SOBSTRIDE = <stride for superobing> ] (1*)     [ FIELDBUFFERSIZE = <fieldbuffersize> ] (1*)     [ INFLATION = <inflation> [ <VALUE>* | PLAIN ] (1*)     [ REGION = <name> <lon1> <lon2> <lat1> <lat2>     ...     [ POINTLOG <lon> <lat> [grid name]]     ...     [ EXITACTION = { BACKTRACE* | SEGFAULT } ]     [ BADBATCHES = <obstype> <max. bias> <max. mad> <min # obs.> ]     ...     [ ZSTATINTS = [<z1> <z2>] ... ]     [ NCFORMAT = { CLASSIC | 64BIT | NETCDF4 } ] (NETCDF4*)     [ NCCOMPRESSION = <compression level> ] (0*)         Notes:     1. { ... | ... | ... } denotes the list of possible choices     2. [ ... ] denotes an optional input     3. ( ... ) is a note     4. * denotes the default value     5. < ... > denotes a description of an entry     6. ... denotes repeating the previous item an arbitrary number of times     7. Depending on the context, some of the entries may be redundant     8. TIME entry is also a flag for geophysical/non-geophysical system,     depending on the presence of "days since"      The TIME entry also serves as a switch between geophysical and non-geophysical systems. For geophysical systems the coordinates are assumed to be geographic longitude and latitude (unless GEOGRAPHIC = 0 is specified for a grid); the time is assumed to be dimensional, in some physical units; and the localisation radii is assumed to be in kilometres. For non-geophysical systems all the above are assumed to be non-dimensional, and the grid is assumed to be on a plane.
@@ 2.4.2 Model parameter file
The model parameter file describes the composition of the state vector by listing the model variables and specifying the associated grids.
@@
Each model variable is described in a block started by the entry for the variable name. The inflation parameters for a variable, if specified, override the common values set in the main parameter file (sec. 2.8.1). The option APPLYLOG makes it possible to conduct assimilation in log space (sec. 2.15), and RANDOMISE – to specify parameters of “forgetting” model for the variable (sec. 2.14).
EnKF-C permits using multiple model grids. Each model variable must be associated with one of the grids defined in the grid parameter file. See examples/4 for an example.
@@ 2.4.3 Grid parameter file
      >./bin/enkf_prep --describe-prm-format grid         Grid parameter file format:         NAME = <name> [ PREP | CALC ]     [ DOMAIN = <domain name> ]     DATA = <data file name>     (either)     [ HTYPE = { rect | curv | unstr | none } ]     XVARNAME = <X variable name>     YVARNAME = <Y variable name>     (if htype = unstr)     [ TRIVARNAME = <triangle vertice IDS variable name> ]     [ TRINEIVARNAME = <triangle neighbour IDS variable name> ]     (end if)     (or)     HGRIDFROM = <grid name>     (end either)     VTYPE = { z | zt | sigma | hybrid | numeric | none }     [ VDIR = { fromsurf* | tosurf } ]     [ GEOGRAPHIC = { 0 | 1* | 2 } ]     (if vtype = z)     ZVARNAME = <Z variable name>     [ ZCVARNAME = <ZC variable name> ]     [ NUMLEVELSVARNAME = <# of levels variable name> ]     [ DEPTHVARNAME = <depth variable name> ]     (else if vtype = sigma)     CVARNAME = <Cs_rho variable name>     [ CCVARNAME = <Cs_w variable name> ]     [ SVARNAME = <s_rho variable name> ] (uniform*)     [ SCVARNAME = <s_w variable name> ] (uniform*)     [ HCVARNAME = <hc variable name> ] (0.0*)     [ DEPTHVARNAME = <depth variable name> ]     [ MASKVARNAME = <land mask variable name> ]     (else if vtype = hybrid)     AVARNAME = <A variable name>     BVARNAME = <B variable name>     [ ACVARNAME = <AC variable name> ]     [ BCVARNAME = <BC variable name> ]     P1VARNAME = <P1 variable name>     P2VARNAME = <P2 variable name>     [ MASKVARNAME = <land mask variable name> ]     (else if vtype = numeric)     ZVARNAME = <Z variable name>     [ ZCVARNAME = <ZC variable name> ]     [ NUMLEVELSVARNAME = <land mask variable name> ]     [ DEPTHVARNAME = <depth variable name> ]     (end if)     [ STRIDE = <stride for ensemble transforms> ] (1*)     [ SOBSTRIDE = <stride for superobing> ] (1*)     [ ZSTATINTS = [<z1> <z2>] ... ]         [ <more of the above blocks> ]
While the code is supposed to automatically identify the types of horizontal grids used, they can also be specified explicitly. The type of the vertical grid has to be specified. If some grid (say, grid A) has the same horizontal grid as another grid (grid B), the code can be notified of this by entering name of grid A in the field HGRIDFROM for grid B (or vice versa). This saves memory, grid initialisation time, and uses transforms calculated for grid A for all variables defined on grid B.
@@
    rectangular quadrilateral grids aligned with geographic or Cartesian coordinates;
@@
The flag GEOGRAPHIC was introduced in v2.8.0. When set to “1” (“yes”, “true”), it communicates to the code that the grid nodes are defined in geographic coordinates. In that case the grid is rendered using 3D Cartesian coordinates; otherwise it is rendered in 2D geographic or Cartesian coordinates, which is somewhat less expensive. The more strict 3D rendering can be essential in polar regions because of the skewness of grid cells in geographic coordinates there, which can cause failures of the mapping algorithm. One can investigate the necessity of setting this flag to “1” by comparing the number of accepted/rejected observations with and without it.
@@
The flag GEOGRAPHIC set to “2” treats grid coordinates similarly to that for “1” but switches to more robust rendering procedures (based on stereographic projections) and is advised to be used e.g. for grids with artificial discontinuities.
For unstructured grids associated with observed variables it is necessary to specify triangulation details: vertex IDs and neighbour triangle IDs. For unstructured grids stride can be equal to 1 only (i.e. ensemble transforms must be calculated in each grid node).
@@
The type of the vertical grid is defined by entry VTYPE in the grid parameter file. EnKF-C supports the following types of vertical grids: “z” ($Z$); “zt”; “sigma” ($\sigma$); “hybrid” ($\sigma-p$), and “numeric”. It is also possible to define a purely horizontal (two-dimensional) grid by defining its vertical type as “none”.
@@
Vertical grid type “numeric” is supposed to handle general situations described by 3D arrays of layer depths. Grids of this type are defined by specifying variable with cell centres depth coordinates (entry ZVARNAME) and (optionally) variable with cell boundaries depth coordinates (entry ZCVARNAME).
@@
The vertical interpolation is performed linearly assuming constant gradients within layers. This means that at the boundary between layers the variables have middle value between those at the adjacent layer centres. This rule applies to vertical grids of all types but “zt”, for which vertical interpolation assumes constant gradients between layer centres.
@@
“Empty” grids have been introduced in version 2.28.0 to simplify estimation of global scalar parameters. In the KF context a model parameter is simply an unobserved model variable. In EnKF-C model variables are defined in the model parameter file; each model variable must be associated with a grid defined in the grid parameter file. “Empty” grids are used to be associated with global model parameters and are defined by specifying their vertical and horizontal types as “none”, e.g.